A biologist finds that 24% of fruit flies have short wings and 47% have eye defects. Assuming these two events are independent, what is the probability that a randomly selected fruit fly has neither short wings nor eye defects?

Answer by stanbon(75887)

A biologist finds that 24% of fruit flies have short wings and 47% have eye defects. Assuming these two events are independent, what is the probability that a randomly selected fruit fly has neither short wings nor eye defects?
-----
P(not short wings) = 1-0.24 = 0.76
-----
P(not eye defects) = 1-0.47 = 0.53
-------
P(not short wings AND not eye defects) = 0.74*0.53 = 0.3922
------------------------
Cheers,
Stan H.
------------------------